/xwiki/resources/icons/silk/feed.png?cache-version=1649904304000 67 posts

LEO center SDI12 subsoil sensor offline

LEO Center SDI12 subsoil sensors when offline after 3/1/2022 at 9pm.

The power supply to these sensors was found to be dead.

It was replaced on 3/14/2022 around 12:40pm.

Since there was no power to the sensors, data at this offline period will not exist.

Higher Infiltration Rate

On February 1, 2022 the standard bi-weekly rain was altered to explore the impact of the developed surface crust at higher infiltration rates.

The standard schedule is: rain from 7:45am to 2:45pm with Multi2 at approximately 13mm/h.

The plan was to cut the rain one hour short and rain the previous hour at double intensity (7:45am-12:45pm with Multi2 at ~13mm/h and 12:45pm-13:45pm with Multi1 and Multi2 at ~26mm/h).

Overland flow started shortly after switching to the higher infiltration rate. Rain was stopped at ~13:03pm. 

As of right now, the rain rate should not exceed 13mm/h.

Short LEO rain

On 3/8/2021, the standard bi-weekly rain from 7:45am to 2:45pm was cut short at 1pm.

This was done to prevent our VFD pumps from running dry as there was insufficient water in the main LEO irrigation tanks.

The auxiliary tanks water was not pumped into the main tanks early enough.

LEO mast kept raised short term

To facilitate various work on the hillslope and due to lack of personnel, the LEO masts have been kept raised since 6/8/2021. This is a short term measure.

Center Bay CSAT3 Removed and Stored

The CSAT 3 instrument was removed by AB on 12/20/20 and stored in a box with the two other CSAT's above the discovery room (in the attic).

 

LEO CNR4 Center Bay reinstalled

LEO-C_10_-2_0_CNR4 was repurposed back in 8/6/2019 10:30am onwards to the West bay for a remote sensing experiment.

See previous note here, LEO CNR4 Center Bay West Sensor Re-purposed

It has been reinstalled back to the original LEO center slope.

*From 8/6/2019 10:30 am to 3/3/2021 10:15 am, LEO-C_10_-2_0_CNR4 does not have any data as it was physically removed.

LEO East SDI12 5TM and MPS2 power supply failure

Due to a 5V power supply failure in the P02 enclosure and the failure was not caught early, LEO East SDI12 5TM and MPS2 are not available.

The power supply failed 10/8/2020 3:00 am to 11/5/2020 10:30am.

Gravity flow experiment

Gravity flow experiment, July 2020

Prior to the experiment the NovaLynx tipping buckets were calibrated (see this post).

The experiment was repeated separately for each hillslope, since the storage tanks cannot hold enough water for a 12h rain with the Multi 1 circuit on all 3 hillslopes simultaneously.

For each hillslope a 1h rain event was performed to calibrate the rainfall distribution followed by a 12h rain event on the following day to reach a gravity flow regime down to a depth of 50cm.

Details on the rainfall calibration: rainfall_calibration.xlsx

Schedule: 

Date Task Comment 
 CENTER 
 2020-07-06 Spanned ropes on Center hillslope to mark cup locations for rainfall calibration. Masts were raised in the morning. Kept raised until after the 12h rain event.
 Ropes spanned upslope-downslope with pre-labeled marks every 2m from 0-30m.
 2020-07-07 Placed cups on Center hillslope and removed ropes. 
 2020-07-08 Issues with Tipping Bucket counts, switched to cumulative counts to reduce impact. See also NovaLynx calibration
 2020-07-09 Rainfall calibration, Multi 1, Center hillslope.
 Collected and weighed cups after 1h of rain (08:30-09:30).
 Details in attached excel file. Air handler off during the rain.
 2020-07-10 12h rain event (8:30-20:30), Multi 1, Center hillslope.  Air handlers were turned off during rain.
Discharge collection for Center -1 (PE102 and NovaLynx) was leaking. Fixed at 19:40.
 Mast lowered after the rain event.
 EAST 
 2020-07-13 Spanned ropes on East hillslope to mark cup locations for rainfall calibration. Masts were raised in the morning. Kept raised until after the 12h rain event.  
 Ropes spanned upslope-downslope with pre-labeled marks every 2m from 0-30m.
 2020-07-14 Placed cups on East hillslope and removed ropes. 
 2020-07-15 Rainfall calibration, Multi 1, East hillslope.
 Collected and weighed cups after 1h of rain (08:30-09:30).
 Details in attached excel file. Air handler off during the rain.
 2020-07-16 12h rain event (8:30-20:30), Multi 1, East hillslope. Air handlers were turned off during rain.
 Mast lowered after the rain event. 
 WEST 
 2020-07-20 Spanned ropes on West hillslope to mark cup locations for rainfall calibration. Masts were raised in the morning. Kept raised until after the 12h rain event.  
 Ropes spanned upslope-downslope with pre-labeled marks every 2m from 0-30m.
 2020-07-21 Placed cups on West hillslope and removed ropes. 
 2020-07-22 Rainfall calibration, Multi 1, West hillslope.
 Collected and weighed cups after 1h of rain (08:30-09:30).
 Details in attached excel file. Air handler off during the rain.
 Gantry malfunction caused an interruption of several hours during the cup collection.
 2020-07-23 12h rain event (8:30-20:30), Multi 1, West hillslope. Air handlers were turned off during rain.
 Mast lowered after the rain event. 

After the experiment, cups, labeled ropes and cardboard transport boxes used for the cup collection during the rainfall calibration were stored in upper habitat office 10.

NovaLynx calibration and switch to cumulative counts

Prior to the gravity flow experiment, the NovaLynx tipping buckets were calibrated in June/July 2020.

For details of the calibration see here: novalynx_calibration_2020.xlsx

During the calibration we noticed that duplicate counts can occur.

To reduce the issue, the counting of the NovaLynx tipping buckets and PE102 flow meters was switched to cumulative counts on 2020-07-08.

LEO Center 14_-3_0 Load cells 3130 channel swap II

LEO Center 14_-3_0 load cells measurement channel was previously switched on 7/3/2019. 

LEO Center Load cells 3130 channel swap

We experienced an error on 3/30/2020 around 2pm where the load cell was reading an extremely high mass than usual.

This was tracked down to water ingressing onto the load cell and causing it to read incorrectly.

We switched the channel again on 6/9/2020.

Currently (August 2020), we are applying a bead of caulking around the LEO hillslope frame above the load cells to prevent water reaching it.

Tags:
Created by Michael Sibayan on 2022/05/03 16:48
  
This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 13.10.5 - Documentation